# Ca2-VDM: Efficient Autoregressive Video Diffusion Model with Causal Generation and Cache Sharing

### v1

@@ Abstract
With the advance of diffusion models, today’s video generation has achieved impressive quality. To extend the generation length and facilitate real-world applications, a majority of video diffusion models (VDMs) generate videos in an autoregressive manner, i.e., generating subsequent clips conditioned on the last frame(s) of the previous clip. However, existing autoregressive VDMs are highly inefficient and redundant: The model must re-compute all the conditional frames that are overlapped between adjacent clips. This issue is exacerbated when the conditional frames are extended autoregressively to provide the model with long-term context. In such cases, the computational demands increase significantly (i.e., with a quadratic complexity w.r.t. the autoregression step). In this paper, we propose Ca2-VDM, an efficient autoregressive VDM with Causal generation and Cache sharing. For causal generation, it introduces unidirectional feature computation, which ensures that the cache of conditional frames can be precomputed in previous autoregression steps and reused in every subsequent step, eliminating redundant computations. For cache sharing, it shares the cache across all denoising steps to avoid the huge cache storage cost. Extensive experiments demonstrated that our Ca2-VDM achieves state-of-the-art quantitative and qualitative video generation results and significantly improves the generation speed. Code is available at https://github.com/Dawn-LX/CausalCache-VDM/
@@ 1 Introduction
Figure 1: (a): Existing autoregressive VDMs with bidirectional generation. The conditional frames can be fixed-length [15, 55] or extendable. (b): Our Ca2-VDM, which uses causal generation to enable KV-cache and introduce cache sharing across all denoising timesteps. Cache writing stands for a partial model forward on the denoised frames (i.e., at timestep $t=0$) until the KV-caches of every layer are computed.
Video diffusion models (VDMs) [12, 32, 21, 24] have made significant advancements by benefiting from the powerful diffusion techniques [16, 37, 38] and prior studies on image generation [33, 29, 4]. In contrast to images, VDMs need to capture interactions across multiple frames and generate all frames simultaneously (e.g., a 16-frame clip). This is usually facilitated by the temporal attention in prevailing UNet- or Transformer-based VDMs [43, 24]. They introduce interdependencies during the bidirectional attention computation. Consequently, the training and inference lengths must be aligned, extremely restricting the flexibility of VDMs in real-world applications such as long-term [15] or live-stream [1] video generation. Meanwhile, simply scaling the clip length at inference time breaks the alignment and leads to poor generation quality (e.g., Figure 1(b) in [30]), unless one undertakes time-consuming retraining or fine-tuning.
To address this issue, an effective and prevalent solution is autoregressive VDMs [2, 15, 21]: They are capable of autoregressively generating subsequent clips conditioned on last frame(s) [^1] Image-to-video models can be considered as a special case. of previous clip, as shown in Figure 1(a). However, the autoregression process of existing VDMs is highly inefficient and redundant: The conditional frames constitute the overlapping frames between adjacent autoregression chunks and they are re-computed at each step. This issue is exacerbated when the conditional frames are extended autoregressively to provide the model with long-term context. In such cases, the model must re-compute all the conditional frames concatenated by the previously generated chunks, with a quadratic computational demand w.r.t. the autoregressive step (cf. Figure 8 in Sec. 4.3).
@@
  • Cache Storage. During inference, the VDM is repeatedly called in the denoising process at each AR step, where each call is taken with a different timestep $t$. All most all Existing VDMs [21, 32] use the same timestep embedding (indexed by $t$) for both conditional and noisy frames. This requires each denoising step to have its own cache, i.e., caching the key/value features for all denoising steps will consume huge GPU memory.
@@
Equipped with causal generation and cache sharing, we propose to store the KV-cache in a queue so that the model can exploit the long-term context while maintaining an affordable computation and storage cost. To support this queue design, the training samples are partially noised to keep clean prefix frames (with random length) as the condition, and the maximum condition length covers the length of KV-cache queue at inference time. Meanwhile, sinusoidal spatial and temporal positional embeddings (i.e., SPEs and TPEs) are added to the frame sequence following Vision Transformer (ViT) [7]. During inference, the TPEs are assigned chunk-by-chunk as the autoregression progresses. To ensure TPEs are correctly assigned when the cumulatively generated video exceeds the training length, we carefully design a cyclic shift mechanism: Cyclic-TPEs 222 Originally, TPEs are re-assigned from scratch at each AR step. However, when KV-cache is enabled, early TPEs have been bound to previous KV-caches. They can not be re-assigned (cf. Figure 4(c) for more details)..
We evaluated our Ca2-VDM on multiple public datasets including MSR-VTT [47], UCF-101 [39], and Sky Timelapse [51] for both text-to-video and video prediction tasks. The results show that our model achieves significant inference speed improvement while maintaining comparable quantitative and qualitative performance as state-of-the-art VDMs. In summary, we make three contributions in this paper: 1) A causal generation structure that allows the intermediate features of conditional frames can be cached and reused in every autoregression step, eliminating the redundant computation. 2) A cache sharing strategy implemented on the KV-cache queue and facilitated by Cyclic-TPEs. It allows the model to acquire extendable context while significantly reducing the storage cost. 3) Our Ca2-VDM achieves comparable performance with SOTA VDMs at a much less computation demand and a high inference speed.
@@ 2 Related Work
Video Diffusion Models (VDMs) have shown impressive generation capabilities, building on the success of latent diffusion models in image generation applications [33, 29, 4]. Some works [22, 19, 17, 54] develop training-free methods for zero-shot video generation based on pretrained image diffusion models (e.g., Stable Diffusion [33]). To leverage video training data and improve the generation quality, many works [9, 12, 43, 32, 6] extend the 2D Unet in text-to-image diffusion models with temporal attention layers or temporal convolution layers. Recent studies [24, 21] also build VDMs based on spatial-temporal Transformers due to their inherent capability of capturing long-term temporal dependencies. We build our Ca2-VDM based on spatial-temporal Transformers following prior structures.
Tuning-free Video Extrapolation. Prior studies have explored autoregressively extrapolating videos using pretrained short video diffusion models without additional finetuning. These methods usually consist of initializing noise sequence based on the DDIM inversion [37, 25] of previously generated frames [28], co-denoising overlapped short clips [42], or iteratively denoising short clips with noise-rescheduling [30]. However, their generation quality is upper-bounded by the pretrained VDMs. Meanwhile, the lack of finetuning also leads to temporal inconsistencies between short clip transitions.
Past-frame Conditioned Video Prediction. To enhance generation quality and temporal consistency, a popular paradigm is training VDMs conditioned on past frames to predict future frames, enabling video extrapolation through autoregressive model calls. Recent works of autoregressive VDMs have studied a variety of design choices for injecting conditional frames, such as adaptive layer normalization [41, 21], cross-attention [53, 21, 15], and explicitly concatenating to the noisy latent along the temporal-axis [13, 21] or channel-axis [5, 10, 50]. Some works [46, 11] also inject conditional frames by adapter-like subnets (e.g., T2I-adapter [26] or ControlNet [52]). In contrast to existing works, our Ca2-VDM avoids the redundant computation of conditional frames by causal generation and cache sharing, and significantly improves the generation speed.
@@ 3 Method
Figure 3: Overview of the Ca2-VDM pipeline. (a): During training, we randomly set $P$ frames clean prefix, and set distinctive timestep embeddings, i.e., tEmb$(0)$ for the clean prefix and $(t)$ for the denoising target. (b): During inference, in each autoregression (AR) step, the model denoises an $l$-frame chunk conditioned on the spatial/temporal KV-caches shared across all timesteps (denoising stage), and then computes the keys/values of denoised chunk to update the KV-caches (cache writing stage). (c): Causal generation block. We further illustrate the details of causal temporal attention with Cyclic-TPEs in Figure 4 and the prefix-enhanced spatial attention in Figure 5.
@@ 3.1 Preliminaries and Problem Formulation
Preliminaries. Diffusion Models [36, 16] are generative models that model a target distribution $\bm{x}_{0}\sim q(\bm{x})$ by learning a denoising process with arbitrary noise levels. To do this, a diffusion process is defined to gradually corrupt $\bm{x}_{0}$ with Gaussian noise. Each diffusion step is $q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I})$, where $t=1,\ldots,T$ and $\beta_{t}\in(0,1)$ is the variance schedule. By applying the reparameterization trick [16], each $\bm{x}_{t}$ can be sampled as $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. Given the diffusion process, a diffusion model is then trained to approximate the reverse process (denoising process). Each denoising step is parameterized as $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\bm{\Sigma}_{\theta}(\bm{x}_{t},t))$, where $\theta$ contains learnable parameters.
Problem Formulation. Following existing mainstream VDMs [12, 21, 24], we develop Ca2-VDM based on latent diffusion models [33] to reduce the modeling complexity of high dimensional visual data. This is achieved by using a pretrained variational autoencoder (VAE) encoder $\mathcal{E}$ to compress $\bm{x}_{0}$ into a lower-dimensional latent representation, i.e., $\bm{z}_{0}=\mathcal{E}(\bm{x}_{0})$. Consequently, the diffusion and denoising processes are implemented in the latent space, formulated as $q(\bm{z}_{t}|\bm{z}_{t-1})$ and $p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t})$, respectively. The denoised latent $\hat{\bm{z}}_{0}$ is decoded back to the pixel space by the pretrained VAE decoder $\mathcal{D}$, i.e., $\hat{\bm{x}}_{0}=\mathcal{D}(\hat{\bm{z}}_{0})$.
In our setting, the model takes as input a VAE encoded latent sequence333 Throughout this paper, we use “$a\!:\!b$” to denote a half-open interval ranging from $a$ (inclusive) to $b$ (exclusive)  $\bm{z}_{0}^{0:L}=[\bm{z}_{0}^{0},\dots,\bm{z}_{0}^{L-1}]\in\mathbb{R}^{L\times H\times W\times C}$ , where $L$ is the number of frames, $H\times W$ is the downsampled resolution, and $C$ is the number of channels. Then, it aims to generate future frames conditioned on past frames, by learning a distribution $p_{\theta}(\bm{z}_{0}^{P:L}|\bm{z}_{0}^{0:P})$. Here the first $P$ prefix frames serve as condition (referred to as clean prefix), and the remaining $L-P$ frames are those to be denoised (referred to as denoising target). The model parameterized by $\theta$ is denoted as $\bm{\epsilon}_{\theta}(\bm{z}_{t}^{0:L},t)$.
@@ 3.2 Causal Generation and Training Objectives
We first introduce the training objectives, followed by an elaboration of the causal generation block (cf. Figure 3(c)). Here we mainly illustrate the causal temporal attention and prefix-enhanced spatial attention layers. As for the visual-text cross attention, it is widely used in VDMs for text-to-video generation [33, 4]. And it is optional for pure video prediction [21]. Since it is not our primary focus, we refer readers to related works [33, 4] for more details.
Training Objectives. Existing diffusion models [16, 27, 29] are trained with the variational lower bound of $\bm{z}_{0}$’s log-likelihood, formulated as $\mathcal{L}_{\text{vlb}}(\theta)=-\log p_{\theta}(\bm{z}_{0}|\bm{z}_{1})+\sum_{t}D_{KL}(q(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0})\|p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t}))$, where $D_{KL}$ is determined by the mean $\bm{\mu}_{\theta}$ and covariance $\bm{\Sigma}_{\theta}$. By re-parameterizing $\bm{\mu}_{\theta}$ as a noise prediction network $\bm{\epsilon}_{\theta}$ and fixing $\bm{\Sigma}_{\theta}$ as a constant variance schedule [16], the model can be trained using a simplified objective function:
@@
In our setting, each sample is partially noised. We randomly keep $P$ consecutive frames uncorrupted as the clean prefix, and the remaining frames are treated as the denoising target, as shown in Figure 3(a). We use different timestep embeddings for the clean prefix (i.e., tEmb$(0)$) and the denoising target (i.e., $(t)$), rather than a unified timestep embedding for the whole video clip as in many existing VDMs [21, 24]. This ensures the cache from the clean prefix can be correctly shared across each denoising timestep $t$ at inference time (since the clean prefix is always assigned with tEmb$(0)$). Consequently, the simplified objective function for our model is
@@
where $[\bm{\cdot},\bm{\cdot}]$ stands for concatenation along the temporal axis, and $\bm{t}$ is the timestep vector with $\bm{t}_{i}=t$ if $i\geq P$ else 00. $\bm{m}\in\{0,1\}^{N}$ is a loss mask to exclude the clean prefix part, i.e., with $\bm{m}_{i}=1$ if $i\geq P$ else 00. In practice, we train the model with learnable covariance $\bm{\Sigma}_{\theta}$ by optimizing a combination of ~$\widetilde{\mathcal{L}}_{\text{simple}}$ and $\mathcal{L}_{\text{vlb}}$ (with the same loss mask) following [27, 29]. More details are left in the Appendix.
Causal Temporal Attention. To introduce the causality, we mask the attention map to force each frame to only attend to its preceding frames, as shown in Figure 4(a). Specifically, the input to each layer is first permuted by treating the spatial resolution $H\times W$ as the batch dimension, and then linearly projected to query, key, and value features as $\bm{Q,K,V}\in\mathbb{R}^{L\times C^{\prime}}$ (for every spatial grid). The causal attention is computed as (we only describe one attention head and omit the diffusion step $t$ for brevity):
@@
where $\bm{M}\in\mathbb{R}^{L\times L}$ is a lower triangular attention mask with $\bm{M}_{i,j}=-\infty$ if $i<j$ else 00.
Prefix-Enhanced Spatial Attention. In analogy to causal temporal attention, integrating the clean prefix and denoising target into one attention sequence helps enhance the guidance of conditional information. Inspired by prior works [18, 32], we do this via spatial-wise concatenation, as shown in Figure 5. Let $\bm{h}_{t}^{0:L}\in\mathbb{R}^{L\times H\times W\times C^{\prime}}$ be the hidden input to each layer, where the number of frames $L$ is treated as batch dimension and $H\times W$ is flattened for attention calculation. We take a sub-prefix of length $P^{\prime}$ and concatenate it to the denoising target. Specifically, for $\bm{h}_{t}^{i}$ from the $i$-th frame, the query is calculated as $\bar{\bm{Q}}(i)=\bm{W}^{Q}\bm{h}_{t}^{i}$. The prefix-enhanced key is calculated as
@@
where $[\bm{\cdot};\bm{\cdot}]$ stands for concatenation along the spatial dimension, and $\bm{h}_{0}^{i}$ is broadcasted by self-repeat $P^{\prime}$ times for every $i<P$ (i.e., the clean prefix part). We do the same operation to obtain the prefix-enhanced value $\bar{\bm{V}}$. Consequently, for every frame, the prefix-enhanced spatial attention is computed as $\text{Attention}(\bar{\bm{Q}},\bar{\bm{K}},\bar{\bm{V}})$ with an attention map of shape $(HW)\times((P^{\prime}+1)HW)$. In practice, we set $P^{\prime}$ relatively small (e.g., $P^{\prime}=3$), as the computational cost scales proportionally with $P^{\prime}$, while adjacent prefix frames tend to exhibit similar appearances. We empirically show that prefix enhancement improves the quality of autoregressive generation (cf. Table 3 in Sec. 4.2).

### v2

@@ Abstract
With the advance of diffusion models, today’s video generation has achieved impressive quality. To extend the generation length and facilitate real-world applications, a majority of video diffusion models (VDMs) generate videos in an autoregressive manner, i.e., generating subsequent clips conditioned on the last frame(s) of the previous clip. However, existing autoregressive VDMs are highly inefficient and redundant: The model must re-compute all the conditional frames that are overlapped between adjacent clips. This issue is exacerbated when the conditional frames are extended autoregressively to provide the model with long-term context. In such cases, the computational demands increase significantly (i.e., with a quadratic complexity w.r.t. the autoregression step). In this paper, we propose Ca2-VDM, an efficient autoregressive VDM with Causal generation and Cache sharing. For causal generation, it introduces unidirectional feature computation, which ensures that the cache of conditional frames can be precomputed in previous autoregression steps and reused in every subsequent step, eliminating redundant computations. For cache sharing, it shares the cache across all denoising steps to avoid the huge cache storage cost. Extensive experiments demonstrated that our Ca2-VDM achieves state-of-the-art quantitative and qualitative video generation results and significantly improves the generation speed. Code is available: https://github.com/Dawn-LX/CausalCache-VDM
@@ 1 Introduction
Figure 1: (a): Existing autoregressive VDMs with bidirectional generation. The conditional frames can be fixed-length (Henschel et al., 2025; Zheng et al., 2024) or extendable. (b): Our Ca2-VDM, which uses causal generation to enable KV-cache and introduce cache sharing across all denoising timesteps. Cache writing stands for a partial model forward on the denoised frames (i.e., at timestep $t=0$) until the KV-caches of every layer are computed.
Video diffusion models (VDMs) (Guo et al., 2024b; Ren et al., 2024; Lu et al., 2024; Ma et al., 2025) have made significant advancements by benefiting from the powerful diffusion techniques (Ho et al., 2020; Song et al., 2021a, b) and prior studies on image generation (Rombach et al., 2022; Peebles & Xie, 2023; Chen et al., 2024a). In contrast to images, VDMs need to capture interactions across multiple frames and generate all frames simultaneously (e.g., a 16-frame clip). This is usually facilitated by the temporal attention in prevailing UNet- or Transformer-based VDMs (Wang et al., 2023b; Ma et al., 2025). They introduce interdependencies during the bidirectional attention computation. Consequently, the training and inference lengths must be aligned, extremely restricting the flexibility of VDMs in real-world applications such as long-term (Henschel et al., 2025) or live-stream (Alonso et al., 2024) video generation. Meanwhile, simply scaling the clip length at inference time breaks the alignment and leads to poor generation quality (e.g., Figure 1(b) in (Qiu et al., 2024)), unless one undertakes time-consuming retraining or fine-tuning.
To address this issue, an effective and prevalent solution is autoregressive VDMs (Blattmann et al., 2023a; Henschel et al., 2025; Lu et al., 2024): They are capable of autoregressively generating subsequent clips conditioned on last frames of previous clip, as shown in Figure 1(a). However, the autoregression process of existing VDMs is highly inefficient and redundant: The conditional frames constitute the overlapping frames between adjacent autoregression chunks and they are re-computed at each step. This issue is exacerbated when the conditional frames are extended autoregressively to provide the model with long-term context. In such cases, the model must re-compute all the conditional frames concatenated by the previously generated chunks, with a quadratic computational demand w.r.t. the autoregressive step (cf. Figure 6 in Sec. 4.3).
@@
  • Cache Storage. During inference, the VDM is repeatedly called in the denoising process at each AR step, where each call is taken with a different timestep $t$. All most all Existing VDMs (Lu et al., 2024; Ren et al., 2024) use the same timestep embedding (indexed by $t$) for both conditional and noisy frames. This requires each denoising step to have its own cache, i.e., caching the key/value features for all denoising steps will consume huge GPU memory.
@@
Equipped with causal generation and cache sharing, we propose to store the KV-cache in a queue so that the model can exploit the long-term context while maintaining an affordable computation and storage cost. To support this queue design, the training samples are partially noised to keep clean prefix frames (with random length) as the condition, and the maximum condition length covers the length of KV-cache queue at inference time. Meanwhile, sinusoidal spatial and temporal positional embeddings (i.e., SPEs and TPEs) are added to the frame sequence following Vision Transformer (ViT) (Dosovitskiy et al., 2020). During inference, the TPEs are assigned chunk-by-chunk as the autoregression progresses. To ensure TPEs are correctly assigned when the cumulatively generated video exceeds the training length, we carefully design a cyclic shift mechanism: Cyclic-TPEs 111 Originally, TPEs are re-assigned from scratch at each AR step. However, when KV-cache is enabled, early TPEs have been bound to previous KV-caches. They can not be re-assigned (cf. Figure 4(c) for more details)..
We evaluated our Ca2-VDM on multiple public datasets including MSR-VTT (Xu et al., 2016), UCF-101 (Soomro et al., 2012), and Sky Timelapse (Zhang et al., 2020) for both text-to-video and video prediction tasks. The results show that our model achieves significant inference speed improvement while maintaining comparable quantitative and qualitative performance as state-of-the-art VDMs. In summary, we make three contributions in this paper: 1) A causal generation structure that allows the intermediate features of conditional frames can be cached and reused in every autoregression step, eliminating the redundant computation. 2) A cache sharing strategy implemented on the KV-cache queue and facilitated by Cyclic-TPEs. It allows the model to acquire extendable context while significantly reducing the storage cost. 3) Our Ca2-VDM achieves comparable performance with SOTA VDMs at a much less computation demand and a high inference speed.
@@ 2 Related Work
Video Diffusion Models (VDMs) have shown impressive generation capabilities, building on the success of latent diffusion models in image generation applications (Rombach et al., 2022; Peebles & Xie, 2023; Chen et al., 2024a). Some works (Lu et al., 2023; Khachatryan et al., 2023; Hong et al., 2023; Zhang et al., 2024) develop training-free methods for zero-shot video generation based on pretrained image diffusion models (e.g., Stable Diffusion (Rombach et al., 2022)). To leverage video training data and improve the generation quality, many works (Ge et al., 2023; Guo et al., 2024b; Wang et al., 2023b; Ren et al., 2024; Dai et al., 2023) extend the 2D Unet in text-to-image diffusion models with temporal attention layers or temporal convolution layers. Recent studies (Ma et al., 2025; Lu et al., 2024) also build VDMs based on spatial-temporal Transformers due to their inherent capability of capturing long-term temporal dependencies. We build our Ca2-VDM based on spatial-temporal Transformers following prior structures.
Tuning-free Video Extrapolation. Prior studies have explored autoregressively extrapolating videos using pretrained short video diffusion models without additional finetuning. These methods usually consist of initializing noise sequence based on the DDIM inversion (Song et al., 2021a; Mokady et al., 2023) of previously generated frames (Oh et al., 2024), co-denoising overlapped short clips (Wang et al., 2023a), or iteratively denoising short clips with noise-rescheduling (Qiu et al., 2024). However, their generation quality is upper-bounded by the pretrained VDMs. Meanwhile, the lack of finetuning also leads to temporal inconsistencies between short clip transitions.
Past-frame Conditioned Video Prediction. To enhance generation quality and temporal consistency, a popular paradigm is training VDMs conditioned on past frames to predict future frames, enabling video extrapolation through autoregressive model calls. Recent works of autoregressive VDMs have studied a variety of design choices for injecting conditional frames, such as adaptive layer normalization (Voleti et al., 2022; Lu et al., 2024), cross-attention (Zhang et al., 2023b; Lu et al., 2024; Henschel et al., 2025), and explicitly concatenating to the noisy latent along the temporal-axis (Harvey et al., 2022; Lu et al., 2024) or channel-axis (Chen et al., 2024b; Girdhar et al., 2024; Zeng et al., 2024). Some works (Weng et al., 2024; Guo et al., 2024a) also inject conditional frames by adapter-like subnets (e.g., T2I-adapter (Mou et al., 2024) or ControlNet (Zhang et al., 2023a)). In contrast to existing works, our Ca2-VDM avoids the redundant computation of conditional frames by causal generation and cache sharing, and significantly improves the generation speed.
@@ 3 Method
Figure 3: Overview of the Ca2-VDM pipeline. (a): During training, we randomly set $P$ frames clean prefix, and set distinctive timestep embeddings, i.e., tEmb$(0)$ for the clean prefix and $(t)$ for the denoising target. (b): During inference, in each autoregression (AR) step, the model denoises an $l$-frame chunk conditioned on the spatial/temporal KV-caches shared across all timesteps (denoising stage), and then computes the keys/values of denoised chunk to update the KV-caches (cache writing stage). (c): Causal generation block. We further illustrate the details of causal temporal attention with Cyclic-TPEs in Figure 4 and the prefix-enhanced spatial attention is left in the Appendix (cf.  Figure 10).
@@ 3.1 Preliminaries and Problem Formulation
Preliminaries. Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are generative models that model a target distribution $\bm{x}_{0}\sim q(\bm{x})$ by learning a denoising process with arbitrary noise levels. To do this, a diffusion process is defined to gradually corrupt $\bm{x}_{0}$ with Gaussian noise. Each diffusion step is $q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I})$, where $t=1,\ldots,T$ and $\beta_{t}\in(0,1)$ is the variance schedule. By applying the reparameterization trick (Ho et al., 2020), each $\bm{x}_{t}$ can be sampled as $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. Given the diffusion process, a diffusion model is then trained to approximate the denoising process. Each denoising step is parameterized as $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\bm{\Sigma}_{\theta}(\bm{x}_{t},t))$, where $\theta$ contains learnable parameters.
Problem Formulation. Following existing mainstream VDMs (Guo et al., 2024b; Lu et al., 2024; Ma et al., 2025), we develop Ca2-VDM based on latent diffusion models (Rombach et al., 2022) to reduce the modeling complexity of high dimensional visual data. This is achieved by using a pretrained variational autoencoder (VAE) encoder $\mathcal{E}$ to compress $\bm{x}_{0}$ into a lower-dimensional latent representation, i.e., $\bm{z}_{0}=\mathcal{E}(\bm{x}_{0})$. Consequently, the diffusion and denoising processes are implemented in the latent space, formulated as $q(\bm{z}_{t}|\bm{z}_{t-1})$ and $p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t})$, respectively. The denoised latent $\hat{\bm{z}}_{0}$ is decoded back to the pixel space by the pretrained VAE decoder $\mathcal{D}$, i.e., $\hat{\bm{x}}_{0}=\mathcal{D}(\hat{\bm{z}}_{0})$.
In our setting, the model takes as input a VAE encoded latent sequence222 Throughout this paper, we use “$a\!:\!b$” to denote a half-open interval ranging from $a$ (inclusive) to $b$ (exclusive)  $\bm{z}_{0}^{0:L}=[\bm{z}_{0}^{0},\dots,\bm{z}_{0}^{L-1}]\in\mathbb{R}^{L\times H\times W\times C}$ , where $L$ is the number of frames, $H\times W$ is the downsampled resolution, and $C$ is the number of channels. Then, it aims to generate future frames conditioned on past frames, by learning a distribution $p_{\theta}(\bm{z}_{0}^{P:L}|\bm{z}_{0}^{0:P})$. Here the first $P$ prefix frames serve as condition (referred to as clean prefix), and the remaining $L-P$ frames are those to be denoised (referred to as denoising target). The model parameterized by $\theta$ is denoted as $\bm{\epsilon}_{\theta}(\bm{z}_{t}^{0:L},t)$.
@@ 3.2 Causal Generation and Training Objectives
We first introduce the training objectives, followed by the causal generation block (cf. Figure 3(c)). Here we focus on the causal temporal attention and prefix-enhanced spatial attention layers. For the visual-text cross attention, it is widely used in VDMs for text-to-video generation (Rombach et al., 2022; Chen et al., 2024a). And it is optional for pure video prediction (Lu et al., 2024). We refer readers to related works (Chen et al., 2024a) for more details.
Training Objectives. Existing diffusion models (Ho et al., 2020; Peebles & Xie, 2023) are trained with the variational lower bound of $\bm{z}_{0}$’s log-likelihood, formulated as $\mathcal{L}_{\text{vlb}}(\theta)=-\log p_{\theta}(\bm{z}_{0}|\bm{z}_{1})+\sum_{t}D_{KL}(q(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0})\|p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t}))$, where $D_{KL}$ is determined by the mean $\bm{\mu}_{\theta}$ and covariance $\bm{\Sigma}_{\theta}$. By re-parameterizing $\bm{\mu}_{\theta}$ as a noise prediction network $\bm{\epsilon}_{\theta}$ and fixing $\bm{\Sigma}_{\theta}$ as a constant variance schedule (Ho et al., 2020), the model can be trained by a simplified objective:
@@
In our setting, each sample is partially noised. We randomly keep $P$ consecutive frames uncorrupted as the clean prefix, and the remaining frames are treated as the denoising target, as shown in Figure 3(a). We use different timestep embeddings for the clean prefix (i.e., tEmb$(0)$) and the denoising target (i.e., $(t)$), rather than a unified timestep embedding for the whole video clip as in many existing VDMs (Lu et al., 2024; Ma et al., 2025). This ensures the cache from the clean prefix can be correctly shared across each denoising timestep $t$ at inference time (since the clean prefix is always assigned with tEmb$(0)$). Consequently, the simplified objective function for our model is
@@
where $[\bm{\cdot},\bm{\cdot}]$ stands for concatenation along the temporal axis, and $\bm{t}$ is the timestep vector with $\bm{t}_{i}=t$ if $i\geq P$ else 00. $\bm{m}\in\{0,1\}^{N}$ is a loss mask to exclude the clean prefix part, i.e., with $\bm{m}_{i}=1$ if $i\geq P$ else 00. In practice, we train the model with learnable covariance $\bm{\Sigma}_{\theta}$ by optimizing a combination of ~$\widetilde{\mathcal{L}}_{\text{simple}}$ and $\mathcal{L}_{\text{vlb}}$ (with the same loss mask) following (Nichol & Dhariwal, 2021; Peebles & Xie, 2023). More details are left in Sec. B.
Causal Temporal Attention. To introduce the causality, we mask the attention map to force each frame to only attend to its preceding frames, as shown in Figure 4(a). Specifically, the input to each layer is first permuted by treating the spatial resolution $H\times W$ as the batch dimension, and then linearly projected to query, key, and value features as $\bm{Q,K,V}\in\mathbb{R}^{L\times C^{\prime}}$ (for every spatial grid). The causal attention is computed as
@@
where $\bm{M}\in\mathbb{R}^{L\times L}$ is a lower triangular attention mask with $\bm{M}_{i,j}=-\infty$ if $i<j$ else 00. Note that we only describe one attention head and omit the diffusion step $t$ for brevity.
Prefix-Enhanced Spatial Attention. In analogy to causal temporal attention, integrating the clean prefix and denoising target into one attention sequence helps enhance the guidance of conditional information. Inspired by prior works (Hu, 2024; Ren et al., 2024), we do this via spatial-wise concatenation (cf.  Figure 10 in the Appendix). Let $\bm{h}_{t}^{0:L}\in\mathbb{R}^{L\times H\times W\times C^{\prime}}$ be the hidden input to each layer, where the number of frames $L$ is treated as batch dimension and $H\times W$ is flattened for attention calculation. We take a sub-prefix of length $P^{\prime}$ and concatenate it to the denoising target. Specifically, for $\bm{h}_{t}^{i}$ from the $i$-th frame, the query is $\bar{\bm{Q}}(i)=\bm{W}^{Q}\bm{h}_{t}^{i}$. The prefix-enhanced key is
@@
where $[\bm{\cdot};\bm{\cdot}]$ stands for concatenation along the spatial dimension, and $\bm{h}_{0}^{i}$ is broadcasted by self-repeat $P^{\prime}$ times for every $i<P$ (i.e., the clean prefix part). We do the same operation to obtain the prefix-enhanced value $\bar{\bm{V}}$. Consequently, for every frame, the prefix-enhanced spatial attention is computed as $\text{Attention}(\bar{\bm{Q}},\bar{\bm{K}},\bar{\bm{V}})$ with an attention map of shape $(HW)\times((P^{\prime}+1)HW)$. In practice, $P^{\prime}$ is relatively small (e.g., $P^{\prime}=3$), as the computational cost scales proportionally with $P^{\prime}$, while adjacent prefix frames tend to exhibit similar appearances. We empirically show that prefix enhancement improves the generation quality (cf. Table 4).